OpenStudy (anonymous):

x + (10/x-2) = x^2+3x/x-2 solve for x...

OpenStudy (anonymous):

\[x+\frac{ 10 }{ x-2 }=\frac{ x^2+3x }{x-2 }\] right?

OpenStudy (anonymous):

\[x=\frac{ x^2+3x }{x-2 }-\frac{ 10 }{x-2 }=\frac{x^2+3x-10 }{ x-2 }\] \[x \left( x-2 \right)=x^2+3x-10\] can you solve it ?

OpenStudy (anonymous):

yes, thanks so much!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

no wait sorry im actually lost .. the way you did it the first response is how its set up on my paper... but did you make up the second response?

OpenStudy (anonymous):

\[x^2-2x=x^2+3x-10,\] \[x^2+3x-x^2+2x=10\] 5x=10 x=10/5=2

OpenStudy (anonymous):

ok thanks